OpenStudy (anonymous):

At 2:00 p.m. two cars start toward each other from towns 240 miles apart. If the rate of one car is 10 mph faster than the other, how fast does each car go if they meet at 5:00 p.m.? If 3x represents the distance that the slower car traveled, then which expression represents the distance the faster car traveled?

OpenStudy (anonymous):

combined speed is 240/3 = 80 mph difference of speeds = 10 x + y = 80 x-y = 10 2x = 90 x = 45 one travelled at 45 mph and other at 35 mph

OpenStudy (anonymous):

distance faster car travelled = 240-3x miles